OpenStudy (anonymous):

I WILL FAN AND MEDAL Factor 9x^6 – 16y^6 completely. First factor each term. 9x^6 = (__)^2 16y^6 = (__)^2 Then use a^2 – b^2 = (a – b)(a + b). Show your work. 9x^6 – 16y^6 =

Directrix (directrix):

First factor each term. 9x^6 = (3x^3)^2 16y^6 = (4y^3)^2 See if you agree with this.

OpenStudy (anonymous):

yes i do @Directrix

Directrix (directrix):

9x^6 – 16y^6 = ( 3x^3 + 4y^3) ( 3x^3 - 4y^3) How about this ^^. It is the difference of two squares just like a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

Yes thats what i got for the final answer. is that correct?

Directrix (directrix):

Based on the instructions, I think that this is all. The sum and difference of two cubes will factor but I don't think you are being asked to do that here.

OpenStudy (anonymous):

okay @Directrix thanks so much

Directrix (directrix):

I am going to add a little work. Just a second.

Directrix (directrix):

9x^6 – 16y^6 = [(3x^3)]^2 - [(4y^3)]^2 --> add this step in your factoring = ( 3x^3 + 4y^3) ( 3x^3 - 4y^3) I think they want it set out like this. @madison.bush That approach explains the reason you were asked to do this: 9x^6 = (__)^2 16y^6 = (__)^2

OpenStudy (anonymous):

@Directrix okay yeah that makes sense. Thank you!

Directrix (directrix):

Add that step. Alrighty, then.

Directrix (directrix):

You're welcome. I'll see you around here on OS.